OpenStudy (acoey7):

A system of equations is shown below: 2x = 5y + 4 3x − 2y = −16 What is the solution to this system of equations?

OpenStudy (adonine):

try solving for x in the first equation and then use substitution.

OpenStudy (adonine):

2x = 5y + 4 solve for the x variable

OpenStudy (adonine):

2x = 2 times x to get x by itself you do the opposite which is division so you dvide by 2

OpenStudy (adonine):

So, x = 5y + 2?

OpenStudy (adonine):

x = 5/2 y + 2 now substitute.

OpenStudy (adonine):

2x = 5y + 4 x = 2.5 + 2 X = 4.5?

OpenStudy (adonine):

(4, 8)